Zelo stock has a beta of 1.23. The risk free rate of return is 2.86 percent and the market rate of return is 11.47 percent. What is the amount of the risk premium on Zelo stock?
Answer #1

Risk premium on Zelo stock:

= Beta×(Market rate of return-Risk free rate)

= 1.23×(11.47%-2.86%)

= 1.23×8.61%

= 10.59%